# Comprehensive Layer-wise Analysis of SSL Models for Audio Deepfake Detection

### v1

@@ Abstract
Yassine El Kheir1,2thanks: Corresponding author: yassine_el_kheir@dfki.de, Youness Samih3, Suraj Maharjan, Tim Polzehl1,2, Sebastian Möller1,2 1German Research Center for Artificial Intelligence (DFKI), Berlin, Germany 2Technical University of Berlin, Berlin, Germany 3IBM Research AI, Abu Dhabi, UAE yassine_el_kheir@dfki.de, tim.polzehl@dfki.de, sebastian.moeller@dfki.de, younes.samih@ibm.com, mhjsuraj@gmail.com
@@ 1 Introduction
While previous studies have highlighted the potential of SSL models in deepfake detection, a thorough layer-wise analysis of SSL models remain largely unexplored. Most current research concentrates on deepfakes created exclusively in English within full-utterance audio scenarios, leaving several important questions unanswered: (1) How effectively do SSL models detect deepfakes across diverse scenarios, including multiple languages, partial, song, and scene deepfakes? (2) How do different layers of SSL models behave in these varying conditions, and is their performance consistent across different setups? (3) Which layers yield the most discriminative features for differentiating real audio from deepfakes?
@@ 4.1 Datasets
Table 1: Mean EER Results Across Datasets Using Full and Partial Transformer Layers of SSL Small Models. # Indicates the Number of Layers Used from the Front-End SSL Models. Best Average EER is Highlighted in yellow.
@@
Table 2: Summary of datasets used in this study. *Evaluated on models trained with LA19; + focused on Track 1.2 (Full Fake) with two eval sets, ADD23.1.2.1 and ADD23.1.2.2. Due to an issue in the test set, we trained on the HAD training set and reported results on the ASVSpoof evaluation set.
EVAL Datasets
@@
Table 3: Mean EER Results Across Datasets Using Full and Partial Transformer Layers of SSL Large Models. # Indicates the Number of Layers Used from the Front-End SSL Models. Best Average EER is Highlighted in yellow.
@@ 5.1 Layer-wise Contribution Analysis
Table 4: Reported average EER across datasets using different Wav2Vec2, Hubert and WavLM SSL models with different back-end layers. AVGin represents the average EER across datasets for each SSL model, calculated by first averaging the EER for different back-ends (FFN and AASIST) within the same reduced-layer model. AVGout indicates the overall average of AVGin, calculated separately for SSL Small and SSL Large models. #L indicates the number of transformer layers.
@@ 6 Small vs Large SSL Models
When comparing Small and Large SSL models across various datasets, no significant performance difference is observed for most datasets, including LA19, LA21, HABLA, PartialSpoof, and SceneFake. However, for more recent datasets, where deepfake models are more advanced, such as DF21, ADD23.1.2.1, and ADD23.1.2.2, larger architectures are essential for capturing subtle deepfake artifacts and improving generalization across languages, as demonstrated with HAD [^6] Models trained on Chinese partial spoof HAD but evaluated on English PartialSpoof.. For instance, Tables 1 and 3 show a 10.95% difference in EER is observed when comparing the best performance of Wav2Vec-Small on ADD23.1.2.1 with Wav2Vec-Large. A 6.2% difference in EER is also noted between WavLM-Small and WavLM-Large on ADD23.1.2.2. On the other hand, Hubert-Large achieves its best performance on CtrSVDD with an EER of 10.53%, lagging behind the best Hubert-Small configuration by 3.19% EER difference.

### v2

@@ Abstract
Yassine El Kheir1,2thanks: Corresponding author: yassine_el_kheir@dfki.de, Youness Samih3, Suraj Maharjan, Tim Polzehl1,2, Sebastian Möller2,1 1German Research Center for Artificial Intelligence (DFKI), Berlin, Germany 2Technical University of Berlin, Berlin, Germany 3IBM Research AI, Abu Dhabi, UAE yassine_el_kheir@dfki.de, tim.polzehl@dfki.de, sebastian.moeller@dfki.de, younes.samih@ibm.com, mhjsuraj@gmail.com
@@ 1 Introduction
While previous studies have highlighted the potential of SSL models in deepfake detection, a thorough layer-wise analysis of SSL models remains largely unexplored. Most current research concentrates on deepfakes created exclusively in English within full-utterance audio scenarios, leaving several important questions unanswered: (1) How effectively do SSL models detect deepfakes across diverse scenarios, including multiple languages, partial, song, and scene deepfakes? (2) How do different layers of SSL models behave in these varying conditions, and is their performance consistent across different setups? (3) Which layers yield the most discriminative features for differentiating real audio from deepfakes?
@@ 4.1 Datasets
Table 1: Mean EER Results Across Datasets Using Full and Partial Transformer Layers of SSL Small Models. # Indicates the Number of Layers Used from the Front-End SSL Models. Best Average EER is Highlighted in yellow. We trained on the HAD training set and reported results on PartialSpoof eval set.
@@
Table 2: Summary of datasets used in this study. *Evaluated on models trained with LA19; + focused on Track 1.2 (Full Fake) with two eval sets, ADD23.1.2.1 and ADD23.1.2.2. We trained on the HAD training set and reported results on PartialSpoof eval set.
Datasets
@@
Table 3: Mean EER Results Across Datasets Using Full and Partial Transformer Layers of SSL Large Models. # Indicates the Number of Layers Used from the Front-End SSL Models. Best Average EER is Highlighted in yellow. We trained on the HAD training set and reported results on PartialSpoof eval set.
@@ 5.1 Layer-wise Contribution Analysis
Table 4: Reported average EER across datasets using different Wav2Vec2, Hubert, and WavLM SSL models with different back-end layers. AVGin represents the average EER across datasets for each SSL model, calculated by first averaging the EER for different back-ends (FFN and AASIST) within the same reduced-layer model. AVGout indicates the overall average of AVGin, calculated separately for SSL Small and SSL Large models. #L indicates the number of transformer layers.
@@ 6 Small vs Large SSL Models
When comparing Small and Large SSL models across various datasets, no significant performance difference is observed for most datasets, including LA19, LA21, HABLA, PartialSpoof, and SceneFake. However, for more recent datasets, where deepfake models are more advanced, such as DF21, ADD23.1.2.1, and ADD23.1.2.2, larger architectures are essential for capturing subtle deepfake artifacts and improving generalization across languages, as demonstrated with HAD [^6] Models trained on HAD training set but evaluated on PartialSpoof eval set.. For instance, Tables 1 and 3 show a 10.95% difference in EER is observed when comparing the best performance of Wav2Vec-Small on ADD23.1.2.1 with Wav2Vec-Large. A 6.2% difference in EER is also noted between WavLM-Small and WavLM-Large on ADD23.1.2.2. On the other hand, Hubert-Large achieves its best performance on CtrSVDD with an EER of 10.53%, lagging behind the best Hubert-Small configuration by 3.19% EER difference.